A particle executes simple harmonic motion with an amplitude of 8.00 cm. At what positions does its speed equal one fifth of its maximum speed?

larger number_____ cm

smaller number _______cm
Answer #1

v = wsqrt(A^2-X^2)

HERE v = 1/5 AW

1/5Aw = wsqrt(A^2-x^2)

squaring on both sides

A^2/25 = A^2-X^2

X^2 = 24A^2/25

X = 1/5AROOT24 = 4.89 cm

x = +4.89 cm

x = -4.89 cm